OpenStudy (anonymous):

@Nnesha The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) -1 -11 0 -1 1 9 g(x) g(x) = 5x + 1 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (anonymous):

x f(X) -1 -11 0 -1 1 9

OpenStudy (anonymous):

the two pints i choose are the 1st and 2nd we are using the formula y2 -y1/x2-x1

Nnesha (nnesha):

i guess that is same question right ??

OpenStudy (anonymous):

so we have -1+11/0-1 1hoch equals -10

OpenStudy (anonymous):

and yes it is :D

OpenStudy (anonymous):

so can you help me .... @Nnesha

Nnesha (nnesha):

nope that is not right

OpenStudy (anonymous):

@Nnesha please

OpenStudy (anonymous):

just this once :(

Nnesha (nnesha):

 that's how you should plug x's like at the top you got positive 11 so 0 -(-1) = ??

Nnesha (nnesha):

one question at at time please

Nnesha (nnesha):

a*

OpenStudy (jdoe0001):

\(\large \begin{array}{rrllll} x&f(x) \\\hline\\ -1&-11\\ 0&-1\\ 1&9 \end{array} ?\)

TheSmartOne (thesmartone):

That is a table of the points.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so LETS start with part a compare w. a sentence

Nnesha (nnesha):

first find slope

OpenStudy (anonymous):

ok

Nnesha (nnesha):

 like at the top you got positive 11 so 0 -(-1) = ?? bec formula to find slope is \[\huge\rm \frac{ y_2 - y_1 }{ x_2 - x_1 }\] there is negative sign at the den. and x_1 is also negative so that suppose to be 0 - (-1)